# SMF: Template-free and Rig-free Animation Transfer using Kinetic Codes

### v1

@@ Abstract
Animation retargeting involves applying a sparse motion description (e.g., 2D/3D keypoint sequences) to a given character mesh to produce a semantically plausible and temporally coherent full-body motion. Existing approaches come with a mix of restrictions – they require annotated training data, assume access to template-based shape priors or artist-designed deformation rigs, suffer from limited generalization to unseen motion and/or shapes, or exhibit motion jitter. We propose Self-supervised Motion Fields (SMF) as a self-supervised framework that can be robustly trained with sparse motion representations, without requiring dataset specific annotations, templates, or rigs. At the heart of our method are Kinetic Codes, a novel autoencoder-based sparse motion encoding, that exposes a semantically rich latent space simplifying large-scale training. Our architecture comprises of dedicated spatial and temporal gradient predictors, which are trained end-to-end. The resultant network, regularized by the Kinetic Codes’s latent space, has good generalization across shapes and motions. We evaluated our method on unseen motion sampled from AMASS, D4D, Mixamo, and raw monocular video for animation transfer on various characters with varying shapes and topology. We report a new SoTA on the AMASS dataset in the context of generalization to unseen motion. (Source code will be released.)
| 1 University College London | 2 Adobe Research |
1 Introduction
@@
Motion brings characters to life. To build any digital twin world, be it scenario planning, games, or movies, there is a fundamental need to add motion to static characters. Manually authoring full-body motion is tedious, error-prone, and requires significant effort from skilled artists. This quickly becomes expensive when scaling to long sequences or animating many different characters. Hence, researchers seek data-driven solutions.
@@
Traditional approaches investigate this problem with explicit priors — statistical body templates (e.g., SMPL [21] for humans, CAFM [39] for animals) to model shape and pose variations or constructing specific character rigs to transfer joint motion to full body motion via skinning weights. These approaches are simple, popular, and fairly robust, but come at the cost of having to first build an expressive shape along with a corresponding pose space.
@@
Automating these workflows with learning-based solutions has gained popularity: learning a space of stick figure character motion (e.g., neural motion field [12]) or phase-based character control [13]); enabling deformation transfer from a source-target pair to a new shape via neural Jacobian fields [2, 26]; or learning skinning weights from a set of annotated rigged characters [41, 32]. These approaches, however, require various levels of intermediate supervision, handle limited either motion or shape variations, and suffer from limited generalization.
@@
Given a coarse motion specification, we transfer animation to a full-character mesh, without access to any rig or morphable template. A natural approach is to treat this as a sequence prediction problem. However, this quickly leads to memory issues as the animations’ length or the meshes’ resolution increases. Increasing the network capacity adversely affects the situation, leading to overfitting as we often have sparse/limited training data. Also, treating the problem at the frame level is efficient, but leads to jittery motion without any temporal coupling.
@@
Inspired by the recent success of latent space diffusion models [34] over pixel space diffusion models, we ask if a similar latent space can be designed for (sparse) motion sequences. To this end, we propose Kinetic Codes, a temporally-informed lightweight motion autoencoder, that we train over a collection of sparse (humanoid) motion sequences across all types of motion. (Since we only rely on keypoints, instead of body meshes/template, we call the latent space kinetic instead of kinematic.) Regularized by this latent space, we train a spatial and a temporal gradient predictor network. We couple the networks through differentiable spacetime integration and supervise the framework, in an end-to-end fashion.
@@
We evaluate SMF for generalization across diverse shapes and unseen motion, and compare against various alternatives. We evaluate our setup on a range of shape (e.g., Mixamo characters, diverse meshes) and motion datasets (e.g., AMASS, D4D, Mixamo motion, monocular video). In summary, we: (i) propose a self-supervised animation transfer framework regularized by kinetic codes, a learned temporally aware latent space; (ii) develop a rig- and template-free animation transfer framework based on simple keypoints as input that is easy to train with sparse supervision and generalizes robustly to new motions and stylized characters; and (iii) report a new SoTA on the AMASS dataset and show realistic motion animation to in-the-wild stylized characters using different 3D as well as 2D coarse space motion specifications.
2 Related Works
@@
Parameterized deformation approaches represent 2D or 3D shapes through a predetermined function of shared parameters and capture deformations as variations of these parameters. A popular example of such models is morphable models [8]. Such techniques encompass cages, explicit formulations [15] or neural approaches [42], blendshapes [17], skinned skeletons [14], Laplacian eigenfunctions [35]. Linking these parameters to the shapes’ surface typically necessitates manual annotation of weights, commonly called weight painting, in 3D authoring tools. Alternatively, with access to sufficient supervision data, data-driven approaches can yield realistic neural rigs, as exemplified by Pinocchio [4], RigNet [41], skinning-based human motion retargeting [23], and skeletal articulations with neural blend shapes [18]. Unsupervised shape and pose disentanglement [44] proposes learning a disentangled latent representation of shape and pose, facilitating motion transfer using shape codes, dependent on registered meshes and maintaining identical connectivity. To plausibly animate these parameterized shapes over time, the parameters should evolve dynamically, weighing the mesh. Although these methodologies require access to body templates and/or rigs, they can still produce jittery results due to loose coupling of individual frame predictions.
Another work, Skeleton-free Pose Transfer for Stylized 3D Characters [20] aims to alleviate the need for rigs by treating the character pose as a set of independent part deformations. By learning the skinning weights and deformations associated with each module, it can match the source pose using linear blending of skinning weights. While it achieves impressive results in pose transfer, being a per-frame method it suffers from artifacts and lacks temporal coherence when applied to animation transfer, as observed in our results. Further, it requires extra inputs, both the source and target shapes in a rest T-pose.
@@
Deep recurrent neural networks are capable of modeling time and shape sequences using LSTMs to predict human joints [10], generate motion in-betweening [11], and to learn a motion field through time [12]. These approaches require access to templates/rigs, and large datasets of joint motion since they are discrete time representations. Qiao et al. [31] utilize mesh convolutions with LSTMs to deform vertices through time; while, Motion Diffusion [33] uses local attention to capture motifs of a single motion and combines it with a diffusion UNet module to produce motion extrapolation and in-betweening. The main challenge is handling long (extrapolation) sequences while still being able to generalize to unseen motion. In this work, we propose using Augmented Neural ODEs [7], operating through temporally-aware kinetic codes, to model time continuously instead of using discretized sequential networks such as LSTMs.
@@
Temporal surface effects can be modeled, physically correctly, by simulating the underlying soft tissues with finite element methods [6, 9]. However, this direct simulation is often slow and requires artists to design the underlying bone and muscle structure [1]. To overcome the stiffness problem and speed up the simulation, reduced-order models have been proposed [25, 28]. When character rigs are available, approaches have been proposed to add soft tissue deformation as an additive per-vertex bump map on top of a primary motion model. Santesteban et al. [36] use this approach, AMASS [22] imparts secondary motion using the blending coefficients of the SMPL shape space [21], while Dyna [29] learns a data-driven model of soft-tissue deformations using a linear PCA subspace. However, these methods require access to primary motion via a skeleton rig and are restricted to humans registered to a canonical template.
PhysDiff [43] uses imitation learning on a physics simulator to train a diffusion network to produce physically-plausible motions for SMPL body shapes. TRAM [40] learns to reconstruct 3D motion in SMPL from in-the-wild videos but requires computing the camera trajectory using SLAM. Instead of limiting to SMPL shapes, SAME [16] presents a skeleton-agnostic framework for animation by embedding the motion space using graph convolutional networks. However, this approach requires supervision using motion pairs, which are difficult to obtain and are limited to biped characters in T-pose. Temporal Residual Jacobian (TRJ) [26] uses NJF and ODE for motion retargeting, and demonstrates good generalization across shapes. However, TRJ requires motion annotation and access to SMPL template during training, has to be retrained for different motion classes, produces jitters due to per-frame prediction, and does not generalize to unseen motion. In Section 4, we compare such methods against SMF.
3 Algorithm
Our goal is to animate unrigged triangulated meshes of 3D characters, conditioned on coarse motion control. These motion control parameters are defined at few keypoints of the body, can be varied in representation (e.g., 2D or 3D), and specify the target pose per frame. We learn to map these coarse keypoint parameters to dense 3D meshes and, as output, generate an animated 3D mesh at every frame described by the input parameters. This nullifies the need for a mesh template or fixed triangulation for both the source motion as well as the target character. Additionally, our method generalizes to unseen motion targets and unseen body shapes, learn from sparse datasets containing a mix of motion examples, and can be applied to long motion sequences.
3.1 Overview
@@
Our inputs are coarse motion parameters, defined as , which characterize the pose required in each frame and the target triangulated mesh defined as . In our experiments, we have tested ranging from 200 (short) to 4000 (long) sequences. We define at fixed keypoints
extracted automatically from a given body using a pretrained pose extractor [38]. Specifically, is a vector at each time step, where is the number of keypoints/joints and is the dimensionality of the chosen input representation. This representation can be 3D keypoint locations on a mesh, 2D keypoints defined on a stick figure frame, or 3D relative Euler angles computed according to the kinematic tree of the chosen pose space. In this work, we focus on results with 2D and 3D keypoints, as we observe their performance to be more robust than 3D relative Euler angles. Our method is self-supervised as we automatically extract the keypoints from mesh sequence.
@@
where describes additional geometric features of the target character such as centroids, normals, and Wave Kernel Signatures [3] computed on the faces of and encoded per-face using a shallow PointNet [30]. These additional features help establish correspondence during inference on unseen in-the-wild shapes. We jointly train the PointNet network along with the other networks in our system. Figure 2 presents an overview of our method.
3.2 Kinetic Codes: Temporally Informed Motion Representation
At the core of our method is the shape deformation module . Empirically, we observed that naively passing the extracted coarse motion parameters to impairs training, and leads to poor generalization and artifacts on unseen motions (see ablation in Section 4). This is unsurprising as the individual does not contain any temporal motion information or context. Therefore, we first embed in the latent space of an autoencoder. By coupling information across time, this representation leads to smoother interpolation, thereby enhancing generalization to unseen motions.
@@ 3.2 Kinetic Codes: Temporally Informed Motion Representation
where , are multi-headed attention encoder and decoder networks, respectively; is the per-frame latent motion representation, referred to as kinetic code, of the same dimensionality as , and are the decoded motion displacements; refers to number of frames. The length of the sequence can vary across motion samples. Note that does see information from surrounding frames.
@@
3.3 3D Shape Posing via Deformation Module
Inputs to our shape deformation module comprise the initial configuration, specifically the shape in its canonical rest pose, the learned motion latents , and the geometric features (derived from ). The module is designed to forecast the pose for each time step . While could be trained to directly forecast the target vertex positions, such an approach results in artifacts, including flipped and folded faces. Following neural gradient space processing [2, 26], we use a simple MLP to predict affine transformation matrices, labeled as Jacobians, at the centroids of the mesh faces, thus encoding relative transforms [37]. Specifically, we predict final vertex positions using predicted Jacobians, integrated through a differentiable Poisson solver to solve a system of linear equations. We supervise this using vertex-to-vertex and Jacobian losses to ensure accuracy in matching ground-truth data. This improves shape consistency and can model various deformations by predicting affine transformations to mesh faces. Concretely, we define,
where is the Jacobian of the initial frame shape and are intermediate residuals predicted by , which are added to to produce the Jacobian for the frame. We found that predicting the Jacobians using the Kinetic codes, coupled with residual connections leads to significantly faster and well-behaved convergence and better generalization. Note that is the identity transformation projected on the local basis of each face of by rotation. All Jacobians are predicted in the central coordinate frame defined by the local bases of the faces of .
We can now directly obtain final vertex positions using the Poisson Solver on the predicted Jacobians. However, as we drift away from the pose of the initial shape or when inferring on unseen motion parameters, accumulated error leads to shape inconsistencies and artifacts. Additionally, utilizes only coarse motion parameters and, as such, is unable to produce temporally coherent and smooth motions on dense meshes for long sequences. Hence, we use an Augmented ODE based formulation, described next, to enable temporally coherent predictions.
3.4 Temporally Coherent Motion Prediction

### v2

# Template-free and Rig-free Animation Transfer using Kinetic Codes

@@ Abstract.
Animation retargetting applies sparse motion description (e.g., keypoint sequences) to a character mesh to produce a semantically plausible and temporally coherent full-body mesh sequence. Existing approaches come with restrictions – they require access to template-based shape priors or artist-designed deformation rigs, suffer from limited generalization to unseen motion and/or shapes, or exhibit motion jitter. We propose Self-supervised Motion Fields (SMF), a self-supervised framework that is trained with only sparse motion representations, without requiring dataset-specific annotations, templates, or rigs. At the heart of our method are Kinetic Codes, a novel autoencoder-based sparse motion encoding, that exposes a semantically rich latent space, simplifying large-scale training. Our architecture comprises dedicated spatial and temporal gradient predictors, which are jointly trained in an end-to-end fashion. The combined network, regularized by the Kinetic Codes’ latent space, has good generalization across both unseen shapes and new motions. We evaluated our method on unseen motion sampled from AMASS, D4D, Mixamo, and raw monocular video for animation transfer on various characters with varying shapes and topology. We report a new SoTA on the AMASS dataset in the context of generalization to unseen motion. Code, weights, and supplementary are available on the project webpage at https://motionfields.github.io/.
1. Introduction
@@
Motion brings characters to life. To build any digital twin world, be it for scenario planning, games, or movies, adding motion to static characters is a fundamental requirement. Manually authoring full-body motion is tedious, error-prone, and requires significant effort from skilled artists. This quickly becomes expensive when scaling to long sequences or animating many different characters. Hence, researchers seek data-driven solutions.
@@
Traditional approaches investigate this problem with explicit priors — statistical body templates (e.g., SMPL (Loper et al., 2015) for humans, CAFM (Sun and Murata, 2020) for animals) to model shape and pose variations or constructing specific character rigs to transfer joint motion to full body motion via skinning weights. These approaches are simple, popular, and robust, but come at the cost of having to first build an expressive shape space along with a corresponding pose space, and are restricted to specific templates.
@@
Automating these workflows with learning-based solutions has gained popularity: learning a space of stick figure character motion (e.g., neural motion field (He et al., 2022)) or phase-based character control (Holden et al., 2017)); enabling deformation transfer from a source-target pair to a new shape via neural Jacobian fields (Aigerman et al., 2022; Muralikrishnan et al., 2024); or learning skinning weights from a set of annotated rigged characters (Xu et al., 2020; Qin et al., 2023). These approaches, however, require various levels of intermediate supervision, are limited in their handling of motion or shape variations, and suffer from limited generalization.
@@
Given a coarse motion specification, we transfer animation to a full-character mesh, without access to any rig or morphable template, at train or test times. A natural approach is to treat this as a sequence prediction problem. However, this quickly leads to memory issues as the animations’ length or the meshes’ resolution increases. Increasing the network capacity adversely affects the situation, leading to overfitting as we often have sparse/limited training data. Also, treating the problem at the frame level is efficient, but leads to jittery motion without any temporal coupling.
@@
Inspired by the recent success of latent space diffusion models (Rombach et al., 2021) over pixel space diffusion models, we ask if a similar latent space can be designed for (sparse) motion sequences. To this end, we propose Kinetic Codes, a temporally-informed lightweight motion autoencoder, that we train over a collection of sparse (humanoid) motion sequences across all types of motion. (Since we only rely on keypoints, instead of body meshes/template, we call the latent space kinetic instead of kinematic.) Regularized by this latent space, we train a spatial and a temporal gradient predictor network. We couple the networks through differentiable spacetime integration and supervise the framework, in an end-to-end fashion.
@@
We evaluate SMF for generalization across diverse shapes and unseen motion, and compare against various alternatives. We evaluate our setup on a range of diverse shapes and motion datasets (e.g., AMASS, D4D, Mixamo, monocular video). In summary, we: (i) propose a self-supervised animation transfer framework regularized by kinetic codes, a learned temporally aware latent space; (ii) develop a rig- and template-free animation transfer framework based on simple keypoints as input that is easy to train with sparse supervision and generalizes robustly to new motions and stylized characters; and (iii) report a new SoTA on the AMASS dataset and show realistic motion transfer to in-the-wild stylized characters using different 3D as well as 2D coarse space motion specifications.
Code, weights, and supplementary are available on the project webpage at https://motionfields.github.io/.
2. Related Works
@@
Parameterized deformation approaches represent 2D or 3D shapes through a predetermined function of shared parameters and capture deformations as variations of these parameters. A popular example of such models is morphable models (Egger et al., 2019). Such techniques encompass cages, explicit formulations (Ju et al., 2005) or neural approaches (Yifan et al., 2020), blendshapes (Lewis et al., 2014), skinned skeletons (Jacobson et al., 2014), Laplacian eigenfunctions (Rong et al., 2008), etc. Linking these parameters to the shapes’ surface typically necessitates manual annotation of weights, commonly called weight painting, in 3D authoring tools. Alternatively, with access to sufficient supervision data, data-driven approaches can yield realistic neural rigs, as exemplified by Pinocchio (Baran and Popović, 2007), RigNet (Xu et al., 2020), skinning-based motion retargeting (Marsot et al., 2023; Aberman et al., 2020; Zhang et al., 2023), and skeletal articulations with neural blend shapes (Li et al., 2021a). Unsupervised shape and pose disentanglement (Zhou et al., 2020) proposes learning a disentangled latent representation of shape and pose, facilitating motion transfer using shape codes, dependent on registered meshes and maintaining identical connectivity. To plausibly animate these parameterized shapes over time, the parameters should evolve dynamically, weighing the mesh. Although these methodologies require access to body templates and/or rigs, they can still produce jittery results due to loose coupling of individual frame predictions.
Notably, Skeleton-free pose transfer (Liao et al., 2022) aims to alleviate the need for rigs by treating the character pose as a set of independent part deformations. By learning the skinning weights and deformations associated with each module, it can match the source pose using linear blending of skinning weights. While it achieves impressive results in pose transfer, being a per-frame method, it suffers from artifacts and lacks temporal coherence when applied to animation transfer, as observed in our results. Furthermore, its reliance on both source and target shapes being provided in a rest T-pose presents a practical limitation, as such canonical poses are often difficult to define for non-humanoids. This dependency challenges its classification as a truly rig-free method.
@@
Deep recurrent neural networks are capable of modeling time and shape sequences using LSTMs to predict human joints (Fragkiadaki et al., 2015), generate motion in-betweening (Harvey et al., 2020), and to learn a motion field through time (He et al., 2022). These approaches require access to templates/rigs, and large datasets of joint motion since they are discrete time representations. Qiao et al. (2018) utilize mesh convolutions with LSTMs to deform vertices through time; while, Motion Diffusion (Raab et al., 2023) uses local attention to capture motifs of a single motion and combines it with a diffusion UNet module to produce motion extrapolation and in-betweening. The main challenge is handling long (extrapolation) sequences while still being able to generalize to unseen motion. In this work, we propose using Augmented Neural ODEs (Dupont et al., 2019), operating through temporally-aware kinetic codes, to model time continuously instead of using discretized sequential networks such as LSTMs.
@@
Temporal surface effects can be modeled, physically correctly, by simulating the underlying soft tissues with finite element methods (Chadwick et al., 1989; Fan et al., 2014). However, this direct simulation is often slow and requires artists to design the underlying bone and muscle structure (Abdrashitov et al., 2021). To overcome the stiffness problem and speed up the simulation, reduced-order models have been proposed (Modi et al., 2020; Park and Hodgins, 2008). When character rigs are available, approaches have been proposed to add soft tissue deformation as an additive per-vertex bump map on top of a primary motion model. Santesteban et al. (2020) use this approach, AMASS (Mahmood et al., 2019) imparts secondary motion using the blending coefficients of the SMPL shape space (Loper et al., 2015), while Dyna (Pons-Moll et al., 2015) learns a data-driven model of soft-tissue deformations using a linear PCA subspace. However, these methods require access to primary motion via a skeleton rig and are restricted to humans registered to a canonical template.
Temporal Residual Jacobian (TRJ) (Muralikrishnan et al., 2024) uses NJF and ODE for motion retargeting, and demonstrates good generalization across shapes. However, TRJ requires motion annotation and also access to SMPL template during training, has to be retrained for different motion classes, produces jitters due to per-frame prediction, and does not generalize to unseen motion. Unlike TRJ, which requires a template to learn motion transfer between different shapes, our self-supervised method reconstructs the original motion on the same shape, eliminating the need for annotated data.
3. Algorithm
Our goal is to animate unrigged triangulated meshes of 3D characters, conditioned on coarse motion signal. These motion control parameters are defined at a few keypoints of the body, can be varied in representation (e.g., 2D or 3D), and specify the target pose per frame. We learn to map these coarse keypoint parameters to dense 3D meshes and, as output, generate an animated 3D mesh at every frame described by the input motion. This nullifies the need for any consistent mesh template or fixed triangulation for both the source motion as well as the target character. Additionally, our method generalizes to unseen motion targets and unseen body shapes, learns from sparse datasets containing a mix of motion examples, and can be applied to long motion sequences.
3.1. Overview
@@
Our inputs are coarse motion parameters, defined as , which characterize the pose required in each frame and the target triangulated mesh defined as . In our experiments, we have tested ranging from 200 (short) to 4000 (long) sequences. We define at fixed keypoints extracted automatically from a given body. Specifically, is a vector at each time step, where is the number of keypoints/joints and is the dimensionality of the chosen input representation. This representation can be 3D keypoint locations on a mesh, 2D keypoints defined on a stick figure frame, or 3D relative Euler angles computed according to the kinematic tree of the chosen pose space. In this work, we focus on results with 2D and 3D keypoints, as we observe their performance to be more robust than 3D relative Euler angles. Our method is self-supervised as we automatically extract the keypoints from mesh sequence.
@@
where describes additional geometric features of the target character such as centroids, normals, and Wave Kernel Signatures (Aubry et al., 2011) computed on the faces of and encoded per-face using a shallow PointNet (Qi et al., 2016). These additional features establish correspondence during inference on unseen in-the-wild shapes. We jointly train the PointNet network and other networks in our system. Figure 2 presents an overview of our method.
Keypoint extraction.
We semi-automatically extract the keypoints . We start with a one-time manual annotation of one human and one animal shape, performed by selecting mesh faces at joint locations. These sparse annotations are then automatically propagated across all characters and motions using point correspondences computed with Diff3F (Dutt et al., 2024). Our framework is
robust to variations between character setups; for example, it bridges the gap even when target keypoint locations (e.g., from Mixamo) differ from the source. This flexibility is possible because our method only requires the number and order of keypoints to be consistent, not their exact spatial positions. This process, where we annotate just two shapes once, allows us to robustly handle diverse character and motion datasets. For extracting 2D keypoints from RGB videos, we use HRNet (Sun et al., 2019), a pretrained pose extractor.
3.2. Kinetic Codes: Temporally-Informed Motion Representation
At the core of our method is the shape deformation module . We observed that naively passing the extracted coarse motion parameters to impairs training, and leads to poor generalization and artifacts on unseen motions (see ablation in Section 4). This is unsurprising as an individual does not contain any temporal motion information or context. Therefore, we first embed in the latent space of an autoencoder. By coupling information across time, this representation leads to smoother interpolation, thereby enhancing generalization to unseen motions.
@@ 3.2. Kinetic Codes: Temporally-Informed Motion Representation
where , are multi-headed attention encoder and decoder networks, respectively; is the per-frame latent motion representation, referred to as kinetic code, of the same dimensionality as , and are the decoded motion displacements; refers to number of frames.
The length of the sequence can vary across motion samples. Note that we use attention on the full sequence as context, and hence a single frame has an understanding of the broader motion sequence.
@@
3.3. 3D Shape Posing via Deformation Module
Inputs to our shape deformation module comprise the initial configuration, specifically the shape in its canonical rest pose, the learned motion latents , and the geometric features (derived from ). The module is designed to forecast the pose for each time step . While could be trained to directly forecast the target vertex positions, such an approach results in artifacts, including flipped and folded faces. Following neural gradient space processing (Aigerman et al., 2022; Muralikrishnan et al., 2024), we use a simple MLP to predict affine transformation matrices, labeled as Jacobians, at the centroids of the mesh faces, thus encoding relative transforms (Sumner and Popović, 2004). Specifically, we predict final vertex positions using predicted Jacobians, integrated through a differentiable Poisson solver to solve a system of linear equations. We supervise this using vertex-to-vertex and Jacobian losses to ensure accuracy in matching ground-truth data; this frees us from requiring additional annotation data. This improves shape consistency and can model various deformations by predicting affine transformations to mesh faces. Concretely, we define,
where is the Jacobian of the initial frame shape and are intermediate residuals predicted by , which are added to to produce the Jacobian for the frame.
We found that predicting the Jacobians using the Kinetic codes, coupled with residual connections leads to significantly faster and well-behaved convergence and better generalization. Note that is the identity transformation projected on the local basis of each face of by rotation. All Jacobians are predicted in the central coordinate frame defined by the local bases of the faces of .
We can now directly obtain final vertex positions using the Poisson Solver on the predicted Jacobians. However, as we drift away from the pose of the initial shape or when inferring on unseen motion parameters, accumulated error leads to shape inconsistencies and artifacts. Additionally, utilizes only coarse motion parameters and, by itself, is unable to produce temporally coherent and smooth motions on dense meshes for long sequences. Hence, we use an Augmented ODE based formulation, described next, to enable temporally coherent predictions.
3.4. Temporally Coherent Motion Prediction